OpenStudy (anonymous):

Integral 1 / ( y√(3+(y^2)))

OpenStudy (anonymous):

3 + y^2 = t use this substitution

OpenStudy (anonymous):

hmm or try this 3 + y^2 = t^2

OpenStudy (anonymous):

2y = 2t*dt/dy dy = t/y*dt

OpenStudy (anonymous):

t ---- dt y*y*(t) \[\frac{1}{y^2}dt\] y^2 = t^2 - 3 \[\frac{1}{t^2 -3}dt\] you can do it now

OpenStudy (anonymous):

thanks